If Bob and Judy combine their savings of $1,800 and $700, respectively, and deposit this amount into an account that pays 9%Please help
Answer #1

Total balance available=(1800+700)=$2500

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=2500*(1+0.09/12)^(12*7)

=2500*1.87320196

=$4683.00(Approx).